OpenStudy (anonymous):

What is the slope of the line that contains the points (-4, -2) and (7, 7)? Enter your answer as a fraction in lowest terms; use a slash ( / ) for the fraction bar.

OpenStudy (anonymous):

slope=(y2 - y1)/(x2 -x1)=(7+2)/(7+4)=9/11